OpenStudy (anonymous):

I can't seem to get a value for the area between these two polar curves: 2+sin(theta) and 3sin(theta) Area = .5 * Integral of (r^2 * dtheta)

OpenStudy (anonymous):

I found the intersection at the top of both curves, so I used pi/2 and 5pi/2 for my bounds.

OpenStudy (anonymous):

I believe it's \[.5\int\limits_{}^{}topfnc^2-bottomfnc^2 d \theta\]

OpenStudy (anonymous):

yes, since there are two equations. 2+sin(theta) is on top, but when I simplify everything and take the integrals, I get 0?!?

OpenStudy (abb0t):

\[\sin(\theta)=\]

OpenStudy (abb0t):

-2/1

OpenStudy (anonymous):

.5 * Integral(4+2sin(theta)-8sin^2(theta) between pi/2 and 5pi/2. Is this right?

OpenStudy (anonymous):

why did you just rearrange that?

OpenStudy (abb0t):

you have two circles. One that starts on the origin to 3 and another -1 to 3.

OpenStudy (anonymous):

I think it should be \[.5\int\limits_{}^{}4+4\sin \theta-8\sin \theta d \theta\] I'm not certain on your bounds though. I'm not saying they're wrong, just that I don't know.

OpenStudy (anonymous):

oh it is 4sin(theta) thanks.

OpenStudy (abb0t):

based on the graph, you should see which is on top and which is on the bottom. You can graph them on wolfram alpha by putting r=2+sin(theta) and r=3sin(theta) and you will see

OpenStudy (anonymous):

to find the bounds, I set the equations = to each other (find the intersection) 3sinx = 2+ sinx 2sinx = 2 sinx = 1 x = pi/2